OpenStudy (anonymous):

A commercial artist is desigining a logo for a local company. The shape of the logo is that of a regular polygon whose interior angles are each 144 degrees. How many sides does the logo have?

OpenStudy (vera_ewing):

144n=(n-2)180 multiply 180 times both the n and the negative 2 144n=180n-360 subtract both sides by 180n so that the variable is on one side only: 144n-180n=180n-360-180n or -36n= -360 divide both sides by -36 n=10 the regular polygon has 10 sides

OpenStudy (anonymous):

Thank you so much!

OpenStudy (vera_ewing):

You're welcome! :-)